# A presentation of the pure cactus group of degree four

### v1

@@ 1. Introduction
The cactus group $J_{n}$ admits a natural surjection onto the symmetric group $S_{n}$ of degree $n$. The kernel of this surjection is called the pure cactus group $PJ_{n}$ of degree $n$. In [9, Theorem 9], Henriques and Kamnitzer showed that $PJ_{n}$ is isomorphic to the fundamental group of the Deligne–Mumford compactification $\overline{M_{0,n+1}}(\mathbb{R})$ of the moduli space of real genus-zero curves with $n+1$ marked points.
@@ 2. Cactus group
For a given integer $n\geq 2$, the cactus group of degree $n$, denoted by $J_{n}$, is defined as the group admitting a presentation with the generators $s_{p,q}$ for $1\leq p<q\leq n$ and the following relations.
@@
Here, $e$ denotes the trivial element of the group and $[p,q]$ denotes the set $\{p,p+1,\dots,q-1,q\}$ of integers for positive integers $p$, $q$ with $p<q$. Throughout the sequel, we denote the generators such as $s_{23}$ (without comma) in stead of $s_{2,3}$ for brevity.
@@
Similar to the braid group, elements of the cactus group can be represented by diagrams of vertical strands on the plane. Some of such diagrams for $J_{4}$ are exhibited in Figure 1.
@@
Due to this diagramatic expression, we see that the cactus group $J_{n}$ admits a natural projection $\pi:J_{n}\to S_{n}$ onto the symmetric group $S_{n}$ of degree $n$, and its kernel is called the pure cactus group of degree $n$, denoted by $PJ_{n}$. See [9, Subsection 3.1] or [7, Section 1] for more details.
We also introduce certain subgroup used in the sections later. For each integer $n\geq 2$ and subset $S\subset[2,n]$, let $J_{n}^{S}$ be the subgroup of $J_{n}$ generated by the elements $s_{p,q}$ for $1\leq p<q\leq n$ and $q-p+1\in S$ and defined by the following relations.
@@
In later sections, we mainly consider the group $J_{4}^{[2,3]}$, which we denote by $J_{4}^{\prime}$ in the rest of paper. It is a subgroup of $J_{4}$ generated by $s_{12},s_{23},s_{34},s_{13},s_{24}$ (except for $s_{14}$), and it has the following presentation.
@@
We denote by $\mathcal{C}$ the Cayley complex of the group $J_{4}^{\prime}$ with respect to the presentation above.
@@ 3. Preliminary
In this section, we prepare some properties of group action for the proof of the Theorem 1.1.
The following two propositions are given in [14]. See [14] for definitions of the terms used.
Proposition 3.1 ([14, Theorem 6.6.13]).
@@
Next we recall the following well-known theorem originated in [13]. Our terminologies follows from [11].
Proposition 3.3 (Poincaré’s polygon theorem, c.f. [11, p223]).
@@
In the following, we give explanations for the terms used in the theorem above.
Let $D$ be a polygon on the hyperbolic plane $\mathbb{H}^{2}$. An identification on $D$ is a map which assigns, to each side $s$, a side $s^{\prime}$ and an isometry $A(s,s^{\prime})$ of $\mathbb{H}^{2}$ so that
@@
    for each side $s$, there exists a neighborhood $V$ of $s$ so that $A(s,s^{\prime})(V\cap D)\cap D=\emptyset$.
@@
$d_{D^{*}}^{*}(x,x^{\prime}):=\inf\sum_{i=e}^{n}d_{\mathbb{H}^{2}}(z_{i},z_{i}^{\prime})$
@@
Let $D$ be a complete polygon with identification. Let $z_{1}$ be some vertex of $D$. There are precisely two sides of $D$ which meet at $z_{1}$; we choose one of these and call it $s_{1}$. There is then a corresponding side $s_{1}^{\prime}$ and a generator $A_{1}=A(s_{1},s_{1}^{\prime})$. Set $z_{2}=A_{1}(z_{1})$ and observe that there is a unique other side $s_{2}$ which has $z_{2}$ as an end point. There is a corresponding side $s_{2}^{\prime}$ and a generator $A_{2}=A(s_{2},s_{2}^{\prime})$. Set $z_{3}=A_{2}(z_{2})$ and observe that there is a unique other side $s_{3}$ which has $z_{3}$ as an end point. In this manner, we get a sequence $\{z_{i}\}$ of vertices, and a sequence $\{A_{i}\}$ of generators. Then there exists periodic sequence of generators $\{A_{1},\cdots A_{n}\}$ such that $A_{n}(\cdots A_{2}(A_{1}(z_{1})))=z_{1}$ which is called a cycle of generators. For a cycle of generators $\{A_{1},\dots,A_{n}\}$, there exists a sequence of vertices $\{z_{1},\dots,z_{n}\}$ in $D$ such that $z_{2}=A_{1}(z_{1})$, …, $z_{n}=A_{n-1}(z_{n-1})$, and $A_{n}(z_{n})=z_{1}$. The sequence of the vertices is called a cycle of vertices. Then, the polygon $D$ is said to satisfy the cycle condition if for each cycle of vertices $\{z_{1},\dots,z_{n}\}$, there exists an integer $\nu$ so that
@@
    $\nu\sum_{i=e}^{m}\alpha(z_{i})=2\pi$,
where $\alpha(z_{i})$ is the angle around $x_{i}$ measured from inside $D$.
Let $D$ be a polygon with identifications. The polygon $D$ is a Poincaré polygon if $D$ is complete and satisfies the cycle condition.
For a cycle of generators $\{A_{1},\dots,A_{n}\}$, by Condition (vii) , the $\nu$ times of the composition $A_{n}\circ\cdots\circ A_{1}$ is the identity on $D$, i.e., we get a relation $(A_{n}\circ\cdots\circ A_{1})^{\nu}=e$ for that cycle. We call such relations the cycle relation.
4. Proof of Theorem 1.1
Before proving Theorem 1.1, we prepare the following key proposition.
@@
As shown in [7, Proposition 2.6], every element of the cactus group has a unique diagram of the normal form. This enable us to enumerate all the elements of $J_{4}^{\prime}$ with bounded word length, equivalently, to determine the local structure, vertices and edges, in $\mathcal{C}$ inside the bounded ball centered at $e$. Note that the vertices of $\mathcal{C}$ are identified with elements in $J_{4}^{\prime}$.
The vertices in $\mathcal{C}$ whose word length is $1$ are just generators;
@@
The following are the vertices whose word length with respect to the generators above is $2$.
@@
As shown Figure 2, There are $25$ edges ($1$-cells) connecting the above vertices (00-cells) in $\mathcal{C}$, and the $2$-cells (quadrangles) containing $e$ on $\mathcal{C}$ are
@@
Therefore, $\mathcal{C}$ is a cell complex such that there are five quadrilateral $2$-cells surrounding each vertices.
We consider the map $f$ corresponding each closed $2$-cell on $\mathcal{C}$ to the square (regular quadrilateral) in $\mathbb{H}^{2}$ whose internal angles are equal to $\frac{2\pi}{5}$. The image of the map $f$ gives a uniform tiling (regular tessellation) of $\mathbb{H}^{2}$ of type $\{4,5\}$. See Figure 3.
@@
We see that this map is an isometry up to scaling $d(f(x),f(y))=Rd(x,y)$ with $R=\cosh^{-1}(\cot^{2}{\frac{\pi}{5}})$ by elementary calculation based on the hyperbolic trigonometry. ∎
In other words, if we set a metric on $\mathcal{C}$ where the length of each edge is the constant $R$, then $\mathcal{C}$ is truly isometric to $\mathbb{H}^{2}$. Therefore, we can apply results for the hyperbolic plane to the Cayley complex $\mathcal{C}$.
Now we consider an action of $PJ_{4}$ on $\mathcal{C}$. The following proposition was essentially obtained in [7].
@@
The map $\Gamma$ induced from $\Gamma_{0}$ defined by the following implies an isometric action of $PJ_{4}$ on $\mathcal{C}$. Note that $\left(\mathcal{C}\right)^{(0)}$ is identified with $J_{4}^{\prime}$.
$\Gamma_{0}:PJ_{4}\times\left(\mathcal{C}\right)^{(0)}$ $\longrightarrow\left(\mathcal{C}\right)^{(0)}$
@@
Moreover, the action $\Gamma$ on $\mathcal{C}$ is free and cocompact.
In the following, for simplicity, we identify $PJ_{4}$ and its image under the action $\Gamma$, and write $g\cdot x$ in stead of $\Gamma(g,x)$ for $g\in PJ_{4}$ and $x\in\mathcal{C}$. In this setting, $PJ_{4}$ is regarded as a subgroup of $\mathrm{Isom}\;\mathcal{C}$.
Now, let $a_{1},\cdots,a_{20}$ be the elements of $J_{4}^{\prime}$ defined as follows.
@@
Then we obtain the next lemma.
Lemma 4.3.
The following elements $g_{1},\dots,g_{10}$ and their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ are contained in $PJ_{4}$.
@@
Furthermore, with respect to the action $\Gamma$ on $\mathcal{C}$ in Proposition 4.2, they are all of the elements of $PJ_{4}$ whose translation lengths are at most 4 and actually equal to $4$. Precisely, if $d(e,g\cdot e)\leq 4$ holds for $g$ in $PJ_{4}$, then $g=g_{i}^{\pm 1}$ and $d(e,g_{i}\cdot e)=4$ for some $1\leq i\leq 10$.
@@
Next we show their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ can be described as in the statement of the lemma.
For $a_{1},\dots,a_{20}$, the following are obtained by straightforward calculations.
@@
Finally, with respect to the action $\Gamma$ on $\mathcal{C}$ in Proposition 4.2, we prove that $g_{1},\dots,g_{10}$ and their inverse are all of the elements of $PJ_{4}$ whose translation lengths are equal to $4$, i.e., if $d(e,g\cdot e)\leq 4$ holds for $g$ in $PJ_{4}$, then $g=g_{i}^{\pm 1}$ and $d(e,g_{i}\cdot e)=4$ for some $1\leq i\leq 10$.
Suppose that an element $g$ of $PJ_{4}$ has translation length at most $4$, i.e. $d(e,g\cdot e)\leq 4$ holds. This means that the element $g\cdot e$ in $J_{4}^{\prime}$ has word length at most 4.
All the elements of $J_{4}^{\prime}$ whose word lengths are $4$ or less can be listed up by brute force enumerations. See Tables 1 and 2. Actually, there are 40 elements of length 3 and 105 elements of length 4.
If $a:=g\cdot e$ in $J_{4}^{\prime}$, then, by the next claim, it lies on the image of $PJ_{4}$ under the map $\mathfrak{p}$, and so $\pi(a)\in\{e,(14)(23)\}$.
@@
Let $\pi$ denotes the natural surjection from $J_{4}$ to $S_{4}$, and $\mathfrak{p}$ the map defined by
$\mathfrak{p}:PJ_{4}$ $\longrightarrow J_{4}^{\prime}$
$g$$\longmapsto\begin{cases}g&g\in J_{4}^{\prime}\\ gs_{14}&g\notin J_{4}^{\prime}.\end{cases}$
Then, the following holds.
$\mathfrak{p}(PJ_{4})=\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$
@@
For any elements $x$ in $\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$, $\pi(x)=e$ or $(14)(23)$. Since $x$ lies on $J_{4}^{\prime}$, $x$ does not contain $s_{14}$ as a word. In the former case, $x$ is in $PJ_{4}$ by definition. Thus, $\mathfrak{p}(x)=x$. In the later case, the following equalities hold.
@@
Therefore, $xs_{14}$ is in $PJ_{4}$ and $xs_{14}$ is not in $J_{4}^{\prime}$. It follows $\mathfrak{p}(xs_{14})=xs_{14}^{2}=x$. Conversely, each element $x$ in $\mathfrak{p}(PJ_{4})$, $x$ lies on $J_{4}^{\prime}$. By the construction of $\mathfrak{p}$, there exists $y$ in $PJ_{4}$ such that $x=y$ or $x=ys_{14}$. It follows that $\pi|_{J_{4}^{\prime}}(x)=e$ or $\pi|_{J_{4}^{\prime}}(x)=(14)(23)$. Therefore, $x$ is in $\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$. ∎
Among the elements in Tables 1 and 2, by direct checking whether the images of the elements under $\pi$ are $e$ or $(14)(23)$, we have only the elements $a_{1},\cdots,a_{20}$ given in the above. (Remark that $a_{16}=s_{34}s_{13}s_{23}s_{24}$ appears as $s_{34}s_{13}s_{24}s_{34}$ in Table 2.)
Consequently, all of elements of $PJ_{4}$ whose translation lengths are at most $4$ are shown to be $g_{1},...,g_{10}$ and their inverse elements $g_{1}^{-1},\dots,g_{10}^{-1}$. ∎
@@
Let $\widetilde{D}$ be the polygon on $\mathcal{C}$ defined by following, which is is visualized in Figure 4.
$\displaystyle\widetilde{D}:=\bigcap_{i=e}^{20}\left\{x\in\mathcal{C}\mid\ d(e,x)\leq d(x,a_{i})\right\}$
@@
From Figure 4, we see the following.
@@
The word length of the vertices $\mathrm{int}\widetilde{D}$ are at most $3$. The vertices on $\partial\widetilde{D}$ whose word length is $3$ are only $s_{13}s_{24}s_{23}$, $s_{23}s_{34}s_{12}$, $s_{24}s_{13}s_{23}$, $s_{34}s_{13}s_{12}$, $s_{12}s_{24}s_{34}$.
We denote the five vertices above by $v_{1}$, …, $v_{5}$, respectively.
We remark that there are ten sides of $\widetilde{D}$ which are not edges in the Cayley complex $\mathcal{C}$. These sides are the diagonals of some 2-cells (quadrangles) in $\mathcal{C}$. The interior angles around $v_{1}$, …, $v_{5}$ are $\frac{2\pi}{5}$. The interior angles around $s_{13}s_{23}$, $s_{13}s_{12}$, $s_{24}s_{34}$, $s_{24}s_{23}$, $s_{34}s_{12}$ are $\frac{4\pi}{5}$. The interior angles around $s_{12}s_{23}$, $s_{13}s_{24}$, $s_{13}s_{34}$, $s_{23}s_{12}$, $s_{23}s_{34}$, $s_{24}s_{12}$, $s_{24}s_{13}$, $s_{34}s_{23}$, $s_{34}s_{13}$, $s_{12}s_{24}$ are $\frac{3\pi}{5}$.
@@
The polygon $\widetilde{D}$ is the Dirichlet polygon with center $e$ for $PJ_{4}$ as a subgroup of $\mathrm{Isom}\;\mathcal{C}$.
@@
Since $PJ_{4}$ is nontrivial, we consider

### v2

@@ 1. Introduction
The cactus group $J_{n}$ admits a natural surjection onto the symmetric group $S_{n}$ of degree $n$. The kernel of this surjection is called the pure cactus group $PJ_{n}$ of degree $n$. In [10, Theorem 9], Henriques and Kamnitzer showed that $PJ_{n}$ is isomorphic to the fundamental group of the Deligne–Mumford compactification $\overline{M_{0,n+1}}(\mathbb{R})$ of the moduli space of real genus-zero curves with $n+1$ marked points.
@@ 2. Cactus group
For any integer $n\geq 2$, we define the cactus group of degree $n$, denoted by $J_{n}$, as the group given by the presentation with generators $s_{p,q}$ for $1\leq p<q\leq n$, subject to the following relations.
@@
Here, $e$ denotes the identity element, and $[p,q]$ denotes the set ${p,p+1,\dots,q-1,q}$ of integers for positive integers $p$ and $q$ with $p<q$. Throughout the sequel, we write generators such as $s_{23}$ (without a comma) instead of $s_{2,3}$ for brevity.
@@
Similar to the braid group, elements of the cactus group can be represented by diagrams of vertical strands on the plane. Some such diagrams for $J_{4}$ are shown in Figure 1.
@@
Due to this diagrammatic expression, we see that the cactus group $J_{n}$ admits a natural projection $\pi:J_{n}\to S_{n}$ onto the symmetric group $S_{n}$ of degree $n$, and its kernel is called the pure cactus group of degree $n$, denoted by $PJ_{n}$. See [10, Subsection 3.1] or [7, Section 1] for more details.
We also introduce a certain subgroup that will be used in later sections. For each integer $n\geq 2$ and subset $S\subset[2,n]$, let $J_{n}^{S}$ be the subgroup of $J_{n}$ generated by the elements $s_{p,q}$ for $1\leq p<q\leq n$ with $q-p+1\in S$, and defined by the following relations.
@@
In later sections, we mainly consider the group $J_{4}^{[2,3]}$, which we denote by $J_{4}^{\prime}$ in the rest of paper. It is a subgroup of $J_{4}$ generated by $s_{12},s_{23},s_{34},s_{13},s_{24}$ (excluding $s_{14}$), and it has the following presentation.
@@
We denote by $\mathcal{G}$ the Cayley graph of the group $J_{4}^{\prime}$, and by $\mathcal{C}$ the Cayley complex of $J_{4}^{\prime}$ with respect to the presentation above. See, for instance, [9] for the definitions of these objects.
@@ 3. Preliminary
In this section, we prepare some properties of group actions for the proof of Theorem 1.1.
The following two propositions are given in [15]. See [15] for the definitions of the terms used.
Proposition 3.1 ([15, Theorem 6.6.13]).
@@
Next, we recall the following well-known theorem, originally from [14]. Our terminology follows that of [12].
Proposition 3.3 (Poincaré’s polygon theorem, c.f. [12, p223]).
@@
In the following, we explain the terms used in the theorem above.
Let $D$ be a polygon in the hyperbolic plane $\mathbb{H}^{2}$. An identification on $D$ is a map that assigns, to each side $s$, a side $s^{\prime}$ and an isometry $A(s,s^{\prime})$ of $\mathbb{H}^{2}$ such that
@@
    for each side $s$, there exists a neighborhood $V$ of $s$ such that $A(s,s^{\prime})(V\cap D)\cap D=\emptyset$.
@@
$d_{D^{*}}^{*}(x,x^{\prime}):=\inf\sum_{i=1}^{n}d_{\mathbb{H}^{2}}(z_{i},z_{i}^{\prime})$
@@
Let $D$ be a complete polygon with identification. Let $z_{1}$ be a vertex of $D$. There are precisely two sides of $D$ meeting at $z_{1}$; we choose one of these and call it $s_{1}$. There is then a corresponding side $s_{1}^{\prime}$ and a generator $A_{1}=A(s_{1},s_{1}^{\prime})$. Set $z_{2}=A_{1}(z_{1})$ and observe that there is a unique other side $s_{2}$ having $z_{2}$ as an endpoint. There is a corresponding side $s_{2}^{\prime}$ and a generator $A_{2}=A(s_{2},s_{2}^{\prime})$. Set $z_{3}=A_{2}(z_{2})$ and observe that there is a unique other side $s_{3}$ having $z_{3}$ as an endpoint. In this way, we obtain a sequence $\{z_{i}\}$ of vertices, and a sequence $\{A_{i}\}$ of generators. Then there exists a periodic sequence of generators $\{A_{1},\dots,A_{n}\}$ such that
$A_{n}(\cdots A_{2}(A_{1}(z_{1})))=z_{1}$
called a cycle of generators. For a cycle of generators $\{A_{1},\dots,A_{n}\}$, there exists a sequence of vertices $\{z_{1},\dots,z_{n}\}$ in $D$ such that
$z_{2}=A_{1}(z_{1}),\ \dots,\ z_{n}=A_{n-1}(z_{n-1}),\text{ and }A_{n}(z_{n})=z_{1}.$
This sequence is called a cycle of vertices. Then, the polygon $D$ is said to satisfy the cycle condition if for each cycle of vertices $\{z_{1},\dots,z_{n}\}$, there exists an integer $\nu$ such that
@@
    $\nu\sum_{i=1}^{n}\alpha(z_{i})=2\pi$,
where $\alpha(z_{i})$ is the interior angle at $z_{i}$ measured inside $D$.
Let $D$ be a polygon with identifications. The polygon $D$ is called a Poincaré polygon if it is complete and satisfies the cycle condition.
For a cycle of generators $\{A_{1},\dots,A_{n}\}$, by Condition (vii), the $\nu$-fold composition
$(A_{n}\circ\cdots\circ A_{1})^{\nu}$
is the identity on $D$, i.e., we have the relation
$(A_{n}\circ\cdots\circ A_{1})^{\nu}=e$
for that cycle. We call such relations the cycle relations.
4. Key propositions
In this section, we present two key propositions, Proposition 4.1 and Proposition 4.2, which are essential for the proof of Theorem 1.1.
@@ 4. Key propositions
As shown in [7, Proposition 2.6], every element of the cactus group has a unique normal form diagram. This enables us to enumerate all the elements of $J_{4}^{\prime}$ with bounded word length, equivalently, to determine the local structure, vertices and edges, in $\mathcal{C}$ inside the bounded ball centered at $e$. Note that the vertices of $\mathcal{C}$ are identified with elements in $J_{4}^{\prime}$.
The vertices in $\mathcal{C}$ whose word length is $1$ are simply the generators;
@@
The following vertices have word length 2 with respect to the generators above.
@@
As shown in Figure 2, there are 25 edges (1-cells) connecting the above vertices (0-cells) in $\mathcal{C}$, and the 2-cells (quadrangles) containing $e$ in $\mathcal{C}$ are
@@
Therefore, $\mathcal{C}$ is a cell complex such that there are five quadrilateral $2$-cells surrounding each vertex.
We consider the map $f$ corresponding to each closed $2$-cell on $\mathcal{C}$ to the square (regular quadrilateral) in $\mathbb{H}^{2}$ whose internal angles are equal to $\frac{2\pi}{5}$. The image of the map $f$ gives a uniform tiling (regular tessellation) of $\mathbb{H}^{2}$ of type $\{4,5\}$. See Figure 3.
@@
We see that this map is an isometry up to scaling, satisfying $d(f(x),f(y))=R\,d(x,y)$ with $R=\cosh^{-1}\left(\cot^{2}{\frac{\pi}{5}}\right)$ by an elementary calculation based on hyperbolic trigonometry. ∎
In other words, if we set a metric on $\mathcal{C}$ by assigning each edge a constant length $R$, then $\mathcal{C}$ is truly isometric to $\mathbb{H}^{2}$. Therefore, we can apply results for the hyperbolic plane directly to the Cayley complex $\mathcal{C}$.
Next, we consider an action of $PJ_{4}$ on the Cayley complex $\mathcal{C}$. The following proposition was essentially obtained in [7].
@@
Let $\Gamma_{0}$ be the action of $PJ_{4}$ on $J_{4}^{\prime}$ defined as follows.
$\Gamma_{0}:PJ_{4}\times J_{4}^{\prime}$ $\longrightarrow J_{4}^{\prime}$
@@
Then $\Gamma_{0}$ induces the maps $\Gamma_{1}$ and $\Gamma$, which give rise to isometric actions of $PJ_{4}$ on $\mathcal{G}$ and $\mathcal{C}$, respectively. Note that $\left(\mathcal{C}\right)^{(0)}=J_{4}^{\prime}$ is identified with $J_{4}^{\prime}$. Moreover, the actions $\Gamma_{1}$ and $\Gamma$ on $\mathcal{G}$ and $\mathcal{C}$ are free and cocompact.
In the following, for simplicity, we identify $PJ_{4}$ with its image under the action $\Gamma$, and write $g\cdot x$ instead of $\Gamma(g,x)$ for $g\in PJ_{4}$ and $x\in\mathcal{C}$. In this setting, $PJ_{4}$ is regarded as a subgroup of $\mathrm{Isom}\;\mathcal{C}$.
5. Proof of Theorem 1.1
In this section, we give our proof of Theorem 1.1. First, we will find a fundamental polygon $\widetilde{D}$ on $\mathcal{C}$, equivalently on $\mathbb{H}^{2}$, with respect to the action $\Gamma$ of $PJ_{4}$, and apply Poincaré’s theorem to obtain a presentation of $PJ_{4}$.
We first prepare the next lemma.
Lemma 5.1.
Let $a_{1},\cdots,a_{20}$ be the elements of $J_{4}^{\prime}$ defined as follows.
@@
Then the following elements $g_{1},\dots,g_{10}$ and their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ are contained in $PJ_{4}$.
@@
Furthermore, with respect to the action $\Gamma_{1}$ on $\mathcal{G}$ in Proposition 4.2, they are all of the elements of $PJ_{4}$ whose translation lengths are at most 4 and actually equal to $4$. Precisely, if $d(e,g\cdot e)\leq 4$ holds for $g$ in $PJ_{4}$, then $g=g_{i}^{\pm 1}$ and $d(e,g_{i}\cdot e)=4$ for some $1\leq i\leq 10$.
@@
Next we show their inverses $g_{1}^{-1},\dots,g_{10}^{-1}$ can be described as in the statement of the lemma. For $a_{1},\dots,a_{20}$, the following are obtained by straightforward calculations.
@@
Finally, with respect to the action $\Gamma_{1}$ on $\mathcal{G}$ in Proposition 4.2, we prove that $g_{1},\dots,g_{10}$ and their inverses are all the elements of $PJ_{4}$ whose translation lengths equal $4$, i.e., if $d(e,g\cdot e)\leq 4$ holds for some $g\in PJ_{4}$, then $g=g_{i}^{\pm 1}$ and $d(e,g_{i}\cdot e)=4$ for some $1\leq i\leq 10$.
Suppose that an element $g$ of $PJ_{4}$ has translation length at most $4$, i.e., $d(e,g\cdot e)\leq 4$ holds. This means that the element $g\cdot e$ in $J_{4}^{\prime}$ has word length at most $4$.
All the elements of $J_{4}^{\prime}$ whose word lengths are at most $4$ can be listed by brute-force enumeration. See Tables 1 and 2. In fact, there are $40$ elements of length $3$ and $105$ elements of length $4$.
If $a:=g\cdot e$ in $J_{4}^{\prime}$, then, by the next claim, it lies in the image of $PJ_{4}$ under the map $\mathfrak{p}$, and so $\pi(a)\in\{e,(14)(23)\}$.
@@
Let $\pi$ denote the natural surjection from $J_{4}$ to $S_{4}$ and $\mathfrak{p}:PJ_{4}\to J_{4}^{\prime}$ the map defined by
$g\longmapsto\begin{cases}g&\text{ if }\ g\in J_{4}^{\prime}\\ gs_{14}&\text{ if }\ g\notin J_{4}^{\prime}.\end{cases}$
Then the following holds:
$\mathfrak{p}(PJ_{4})=\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\}).$
@@
For any element $x$ in $\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$, we have $\pi(x)=e$ or $(14)(23)$. Since $x$ lies in $J_{4}^{\prime}$, it does not contain $s_{14}$ as a word. In the former case, $x$ is in $PJ_{4}$ by definition, and thus $\mathfrak{p}(x)=x$. In the latter case, the following equalities hold.
@@
Thus, $xs_{14}$ is in $PJ_{4}$ and $xs_{14}$ is not in $J_{4}^{\prime}$. It follows that $\mathfrak{p}(xs_{14})=xs_{14}^{2}=x$.
Conversely, for each element $x$ in $\mathfrak{p}(PJ_{4})$, we have $x\in J_{4}^{\prime}$. By the construction of $\mathfrak{p}$, there exists $y\in PJ_{4}$ such that $x=y$ or $x=ys_{14}$. It follows that
$\pi|_{J_{4}^{\prime}}(x)=e\quad\text{or}\quad\pi|_{J_{4}^{\prime}}(x)=(14)(23).$
Therefore, $x\in\left(\pi|_{J_{4}^{\prime}}\right)^{-1}(\{e,(14)(23)\})$. ∎
Among the elements in Tables 1 and 2, by directly checking whether the images of the elements under $\pi$ are $e$ or $(14)(23)$, we obtain only the elements $a_{1},\cdots,a_{20}$ as given above. (Remark that $a_{16}=s_{34}s_{13}s_{23}s_{24}$ appears as $s_{34}s_{13}s_{24}s_{34}$ in Table 2.)
Consequently, all elements of $PJ_{4}$ whose translation lengths are at most $4$ are shown to be $g_{1},\dots,g_{10}$ and their inverse elements $g_{1}^{-1},\dots,g_{10}^{-1}$. ∎
@@
Let $\widetilde{D}$ be the polygon on $\mathcal{C}$ defined as follows, which is visualized in Figure 4.
$\displaystyle\widetilde{D}:=\bigcap_{i=1}^{20}\left\{x\in\mathcal{C}\mid\ d(e,x)\leq d(x,a_{i})\right\}$
@@
From Figure 4, the following can be verified.
@@
With respect to the action $\Gamma_{1}$ on $\mathcal{G}$, the word lengths of the vertices in $\mathrm{int}\,\widetilde{D}$ are at most $3$. The vertices on $\partial\widetilde{D}$ whose word length is $3$ are only $s_{13}s_{24}s_{23}$, $s_{23}s_{34}s_{12}$, $s_{24}s_{13}s_{23}$, $s_{34}s_{13}s_{12}$, $s_{12}s_{24}s_{34}$. ∎
We denote the five vertices above by $v_{1}$, $\dots$, $v_{5}$, respectively. We remark that there are ten sides of $\widetilde{D}$ which are not edges in the Cayley complex $\mathcal{C}$. These sides are the diagonals of some 2-cells (quadrangles) in $\mathcal{C}$. The interior angles at $v_{1}$, $\dots$, $v_{5}$ are $\frac{2\pi}{5}$. The interior angles at the vertices $s_{13}s_{23}$, $s_{13}s_{12}$, $s_{24}s_{34}$, $s_{24}s_{23}$, $s_{34}s_{12}$ on $\partial\widetilde{D}$ are $\frac{4\pi}{5}$. The interior angles at $s_{12}s_{23}$, $s_{13}s_{24}$, $s_{13}s_{34}$, $s_{23}s_{12}$, $s_{23}s_{34}$, $s_{24}s_{12}$, $s_{24}s_{13}$, $s_{34}s_{23}$, $s_{34}s_{13}$, $s_{12}s_{24}$ on $\partial\widetilde{D}$ are $\frac{3\pi}{5}$.
@@
The subset $\widetilde{D}_{1}:=\mathrm{int}(\widetilde{D}\cap\mathcal{G})$ of $\mathcal{G}$ is the Dirichlet domain with center $e$ in $\mathcal{G}$ with respect to the action $\Gamma_{1}$ of $PJ_{4}$.
@@
Since $PJ_{4}$ is nontrivial, by definition, we consider